OpenStudy (anonymous):

how to find the square root of 30

OpenStudy (anonymous):

find the factors of 30 first.

OpenStudy (anonymous):

you might use calculator! otherwise find the square roots that are near to thirty. those are going to be 25 and 36.

OpenStudy (anonymous):

\[\sqrt{2*3*5}=\sqrt{30}\]

OpenStudy (anonymous):

so is that my answer?

OpenStudy (anonymous):

i don,t understand

OpenStudy (anonymous):

yes friend its your answer.

OpenStudy (anonymous):

but i thought it would have to be one number

OpenStudy (phi):

you can use the google search window sqrt(30)= if you want a decimal number. But there is no exact answer except \( \sqrt{30} \)

OpenStudy (anonymous):

\[\sqrt{30}=5.477 its also your answer.\]